Improper Integral (Integrals with Infinite Limits)


∫dx/(x^2-2x-3) from 0 to 4


1
Expert's answer
2021-05-07T14:16:03-0400

"∫_0^4\dfrac{dx}{(x^2-2x-3)}"


"I =∫_0^4\dfrac{dx}{(x^2-2x-3)}"


We can factorise the integrand and decompose into partial fraction as follows:

"\dfrac{1}{(x^2-2x-3)}\equiv\dfrac{1}{(x-3)(x+1)}"


Thus,

"I =∫_0^4\dfrac{dx}{(x^2-2x-3)}= ∫_0^4\dfrac{\frac{1}{4}}{(x-3)}+\dfrac{\frac{-1}{4}}{(x+1)}"


"I =\dfrac{1}{4}∫_0^4\dfrac{1}{(x-3)}- \dfrac{1}{(x+1)}"


"I = \dfrac{1}{4} \ [ ln|x-3| - ln |x+1|]_0^4"


"I =\dfrac14 \ {(ln1-ln5)-(ln3-ln1)}"


"I = \dfrac{1}{4 }\ (0-ln5-ln3-0)"


"I= -\dfrac{1}{4 }\ (ln5+ln3)"